OpenStudy (anonymous):

the sum of 3 consecutive odd integers is -15 what are the numbers?

OpenStudy (lasttccasey):

Just using process of elimination, I came up with: -3, -5, -7 -3+-5+-7=-15

OpenStudy (lasttccasey):

No, it has to be consecutive.

OpenStudy (anonymous):

omg yea thanks

OpenStudy (anonymous):

three consecutive odd numbers: a, a+2, a+4, so a +a+2 + a+4 = -15 = 3 a + 6 3a = -21, a = -7, check: -7, -5, -3 do add to -15. These are consecutive ODD numbers. @lasttccasey was right.